wolvez5002:

deans company offers a reimbursement package of 0.59

wolvez5002:

c=0.59+275x c=0.59x+275 c=59x+275 c=59x+275

wolvez5002:

the question is deans company offers a reimbursement of 0.59 per mile plus 275 a year maintance. if x represents the number of miles which equation models c , the total amount of reimbursement the company offers

Mercury:

old question but will respond so this can be closed (reimbursement per mile)(number of miles) + (yearly maintenance fee) fill in the appropriate values from the problem